### v2

# MIM-Refiner: A Contrastive Learning Boost from Intermediate Pre-Trained Representations

Abstract
We introduce MIM (Masked Image Modeling)-Refiner, a contrastive learning boost for pre-trained MIM models. MIM-Refiner is motivated by the insight that strong representations within MIM models generally reside in intermediate layers. Accordingly, MIM-Refiner leverages multiple contrastive heads that are connected to different intermediate layers. In each head, a modified nearest neighbor objective constructs semantic clusters that capture semantic information which improves performance on downstream tasks, including off-the-shelf and fine-tuning settings.
The refinement process is short and simple – yet highly effective. Within a few epochs, we refine the features of MIM models from subpar to state-of-the-art, off-the-shelf features. Refining a ViT-H, pre-trained with data2vec 2.0 on ImageNet-1K, sets a new state-of-the-art in linear probing (84.7%) and low-shot classification among models that are pre-trained on ImageNet-1K. At ImageNet-1K 1-shot classification, MIM-Refiner advances the state-of-the-art to 64.2%, outperforming larger models that were trained on up to 2000 times more data such as DINOv2-g, OpenCLIP-G and MAWS-6.5B. https://ml-jku.github.io/MIM-Refiner
1 Introduction
Figure 1: Linear probing state-of-the-art on ImageNet-1K over the last four years.
Self-supervised learning (SSL) has attracted considerable attention, owing to its data efficiency and generalization ability [49]. SSL leverages pre-taining tasks and creates intricate input representations without the need for explicit supervision via expensive annotations. In computer vision, Masked Image Modeling (MIM) [12, 69, 58] has emerged as one of the prevalent SSL pre-training paradigms, enabling an efficient pretraining of large models on unlabeled data by reconstructing masked parts of the input images.
The success of MIM is driven by methods like Masked Autoencoder (MAE) [29], data2vec [7, 6], and others [8, 75]. For example, MAE opens the door for sparse pre-training of Vision Transformers (ViTs) [22] by masking large parts of the image and not processing the masked areas. The computational efficiency, coupled with the data efficiency of a generative reconstruction task [76, 25, 64] fosters the scaling to larger architectures on datasets of limited size. However, MIM models tend to spread their attention across the whole image [70]. When adapting to downstream tasks, a sufficient amount of labels is required to rewire the attention to focus on important regions in the image. In the absence thereof, MIM models perform poorly. In contrast, for example, Instance Discrimination (ID) [30, 14, 10, 28] methods implicitly focus on objects and form semantic clusters in the latent space [11], which eases adaption to downstream tasks in the absence of vast amounts of labels. In summary, the most important desiderata for efficient SSL pre-training methods in computer vision are rich representations of the input – ideally in the form of semantic clusters in the latent space – alongside efficiency in both compute and data, and, most notably, favorable scaling to larger architectures.
Figure 2: Left: Representation quality of SSL methods evaluated via linear probing. MIM-Refiner advances state-of-the-art among models pre-trained on ImageNet-1K in low- and high-compute regimes. Right: MIM-Refiner advances state-of-the-art in the extreme setting of 1-shot classification despite being trained on orders of magnitude less data. Size of dots corresponds to model FLOPS.
In this work, we first analyze pre-trained MIM models, where we find that MIM models have different types of blocks: those that mainly improve the pre-training objective and others that are responsible for abstraction within the MIM encoder. The origin of this behavior can be traced back to the fact that MIM architectures usually comprise a large ViT encoder together with a very light-weight decoder. For larger models, the light-weight decoder reaches a point, where it cannot further improve the pre-training objective on its own and passes part of the reconstruction task back to the last encoder blocks. Consequently, the feature quality for downstream tasks of the later blocks degrades, and, somewhat unusual, the representation quality peaks in the middle blocks of the encoder.
Based on these insights, we introduce MIM-Refiner, a simple – yet highly effective – sequential refinement approach tailored to MIM models. MIM-Refiner applies an ensemble of ID heads that enforce semantic clusters via an ID objective. Most importantly, those ID heads are attached to intermediate blocks of the encoder including those that exhibit peak representation quality, instead of only a single ID head attached to the last block.
We introduce Nearest Neighbor Alignment (NNA) as a new ID variant. Given a batch of samples, NNA aligns each sample with its NN and repels each sample from all other samples in the batch.
Experimentally, we show that within few epochs, MIM-Refiner refines the features of MIM models from subpar to state-of-the-art, off-the-shelf features, see Figure 2. For example, a refined ViT-H, pre-trained with data2vec 2.0 [6] on ImageNet-1K [20], achieves new state-of-the-art results in linear probing (84.7%, Figure 1), $k$-NN accuracy (82.3%), $k$-means cluster accuracy (67.3%) and low-shot classification among models pre-trained on ImageNet-1K. In ImageNet-1K 1-shot classification it sets a new state-of-the-art of 64.2%, outperforming larger models that were trained on up to 2000x more data such as DINOv2-g, OpenCLIP-G and MAWS-6.5B.
Our contributions can be summarized as follows:
  1. 1.
    We analyze MIM models, and find that MIM models have different types of blocks: those that mainly improve the pre-training objective and others that are responsible for abstraction within the MIM encoder.
  2. 2.
    We introduce MIM-Refiner, a sequential approach to refine the representation of a pre-trained MIM model to form semantic clusters via an ID objective. Motivated by the findings in (1), MIM-Refiner is designed to exploit the intermediate representations via an ensemble of ID heads attached to multiple encoder blocks.
  3. 3.
    We introduce Nearest Neighbor Alignment (NNA), an ID objective that aligns each sample with its nearest neighbor while repelling other samples in a batch.
2 Challenges of Masked Image Modeling
(a) Encoder-decoder Parameter Count
(b) ImageNet-1K $k$-NN Accuracy per Block
(c) ImageNet-1K Low-shot Accuracy.
(d) Relative Improvements per Block.
Figure 3: (a) MIM models are asymmetrically designed where the encoder has most of the parameters, as indicated by the percentages in the bars. (b) The representation quality of the encoder (measured by $k$-NN accuracy) peaks in the middle blocks before degrading when the later blocks take over parts of the decoder’s task. Various MIM models show such a pattern (see Appendix C.1) (c) The decline in representation quality in later blocks primarily contributes to the degradation in downstream performance. ID methods (represented by iBOT [81]) and our MAE-Refined do not suffer from this issue. (d) Correlation of the relative improvement of the reconstruction loss and the relative improvement in the $k$-NN accuracy per block. The relative $k$-NN accuracy improvement is obtained from the $k$-NN accuracy shown in (b). Similarly, the relative improvement of the reconstruction loss is obtained from Figure 8 in the appendix. Middle blocks form abstract representations (large improvements in the $k$-NN accuracy, almost no improvement in reconstruction loss), later blocks take over parts of the reconstruction task (decrease in the $k$-NN accuracy, large improvement in the reconstruction loss). Additional details can be found in Appendix E.2.
MIM models, such as MAE [29] and data2vec 2.0 [6], enable an efficient pre-training of large models. In terms of architecture, the encoder and decoder are intentionally designed asymmetrically. The encoder, on the one hand, is a large ViT, where discarding 75% of the input patches drastically reduces the cost of a forward pass. The decoder, on the other hand, operates on the full sequence length – by concatenating mask tokens to the encoded visible patches – and, thus, is typically very lightweight to compensate for the increased number of tokens (Figure 3(a)). As models increase in size, the decoder eventually reaches a point where it cannot further improve the pre-training objective on its own. Consequently, it begins to delegate a portion of the reconstruction task back to the last encoder blocks. This transfer adversely affects the feature quality for downstream tasks associated with those blocks (Figure 3(b)), especially when only few labels are available (Figure 3(c)). We observe this phenomenon by correlating the relative improvement in the reconstruction loss vs the $k$-NN accuracy (Figure 3(d)). Roughly speaking, the blocks of the encoder operate in three different regimes:
  1. 1.
    In early ViT blocks, general purpose features are learned, which improve the reconstruction loss and the $k$-NN accuracy simultaneously.
  2. 2.
    In middle ViT blocks, abstractions are formed. The reconstruction loss improves only slightly, while the $k$-NN accuracy improves drastically.
  3. 3.
    In late ViT blocks, features are prepared for the reconstruction task. The reconstruction loss improves at a faster rate, while the $k$-NN accuracy decreases.
When naïvely using the features of the last encoder block, those features are suited for reconstruction, but not particularly for downstream tasks. If lots of labels are available, this can be compensated by fine-tuning the last encoder blocks on the downstream task. However, if not enough labels are available, or the last encoder blocks are not fine-tuned, downstream performance suffers.
One would think that simply using a larger decoder solves these problems. However, there are multiple problems with this solution. (i) The decoder commonly operates on the full sequence length, making it costly to increase its size. (ii) Scaling the decoder can decrease performance as shown, for example, in MAE [29] (Table 1). (iii) Models that can use a deeper decoder (such as CrossMAE [27]) also show degrading representation quality in later blocks, as shown in Appendix Figure 6(c).
3 Method
We propose MIM-Refiner, a novel approach aimed at improving downstream performance by refining the later blocks of a pre-trained MIM model. MIM-Refiner leverages the abstract intermediate representations with an ensemble of Instance Discrimination (ID) heads, which are attached to multiple blocks towards the end of the encoder, as visualized on the left side of Figure 4. The resulting experimental improvements in various downstream tasks are discussed in Section 5.
Figure 4: Left: Comparison of different pre-training schemes. ID uses a single ID head, whereas MIM models use a light-weight decoder to train an encoder. MIM-Refiner attaches multiple ID heads to the later third of the blocks of a pre-trained MIM encoder. Right: Nearest Neighbor Alignment (NNA). An anchor sample is attracted by its NN and simultaneously repelled from other samples in the batch. The NN is retrieved from a first-in-first-out queue of samples from previous batches.
Inspired by NN based contrastive learning [24, 4], we propose Nearest Neighbor Alignment (NNA). NN contrastive objectives introduce an inter-sample correlation by retrieving NNs of samples in a batch and subsequently applying an objective between the samples and their NNs. In practice, the NN retrieval is typically done by tracking samples from previous iterations in a first-in-first-out queue [24, 30]. Therefore, the samples in the queue are from a previous model state which creates a tradeoff between the benefit of the NN-swap and the worse signal from the out-of-sync samples. We argue that the NN-swap does not offer a benefit for negative samples, since there are already different images in every batch, and instead, degrades the signal from the contrastive objective. We therefore propose to use the NN only for the alignment of positive samples, as visualized on the right side of Figure 4. Omitting the NN-swap for the negatives gives a small but consistent improvement as ablated in Table 4(e). NNA is a variant of NNCLR and we visualize their difference in Figure 10.
Formally, given a batch of features $\mathcal{Z}=\{z_{1},\dots,z_{N}\}$ and a queue $\mathcal{Q}$ the NNA objective is:
$\mathcal{L}^{\text{NNA}}_{i}$ $=-\log\frac{\exp(\text{NN}(z_{i},\mathcal{Q})\cdot z_{i}/\tau)}{\exp(\text{NN}(z_{i},\mathcal{Q})\cdot z_{i}/\tau)+\sum\limits^{N}_{j=1}\exp(\text{SG}(z_{j})\cdot z_{i}/\tau)[i\neq j]}$ (1)
$\text{NN}(z_{i},\mathcal{Q})$ $=\underset{q\in\mathcal{Q}}{\text{argmax}}(z_{i}\cdot q)$ (2)
where $z_{i}$ is the anchor, $\text{NN}(z_{i},\mathcal{Q})$ is the positive, $z_{j}$ are the negatives, $\tau$ is the temperature, and SG is the stop-gradient operation. $[i\neq j]$ denotes the Iverson bracket that evaluates to $1$ if $i\neq j$ and to 00 if $i=j$. All vectors are assumed to be normalized to length $1$.
4 Related Work
4.1 Pre-training in Computer Vision
Following the success of generative pre-training of transformers [68] in language modeling [21, 61], similar directions were explored in computer vision [22, 75, 72, 13]. With the introduction of Vision Transformers [22], large Masked Image Modeling (MIM) models could be efficiently trained [29, 6, 64] by using the ability of transformers to effortlessly process sparse input by dropping masked patch tokens in the input and subsequently reconstructing the masked parts. In order to optimize the MIM pre-training objective, models have to infer the missing regions by efficiently encoding foreground and background alike which leads to a rich representation.
Building on rich features learned by MIM models has been explored in various ways. MAWS [64] first pre-trains an MAE, followed by weakly supervised training on a billion-scale dataset using billion-scale models. SemiViT [9] uses a pre-trained MAE as a starting point for semi-supervised learning. Segment Anything [43] use MAE as basis for a segmentation foundation model. MIM-Refiner also builds on the rich features of MIM models and refines them with a ID objective to ease adaption to downstream tasks.
Instance Discrimination (ID) is another branch of self-supervised learning that uses augmentations to create multiple views of the same sample where the task is then to find matching pairs within the views from all samples within a batch [14, 30, 23, 73] or align features of one view with the features from another [28, 11]. We use the terminology that views from the same sample are “positive pairs” and views of different samples are “negative pairs”. When describing a single view of a sample, it is called the “anchor” to which all other views in a batch are either “positives” or “negatives”. NN-based ID [24, 4] extends this setting to use NNs in various ways to create views or otherwise augment samples during training. MIM-Refiner introduces Nearest Neighbor Alignment, which is a modification of previous NN-based ID methods to use the NN only for the alignment part, i.e. pulling the anchor closer to the NN of its positives while pushing the anchor away from its negatives.
4.2 Combining MIM and ID
Adding a MIM to ID methods has emerged as a powerful pre-training scheme. However, in contrast to ID models, the MIM objective in the end-to-end training either uses significantly lower masking ratios with mask tokens getting processed by the encoder [81, 56], or requires a target encoder to encode the unmasked image [37, 2]. Both drastically increase the computational requirements as the encoder operates on the full sequence length. Consequently, these models either require copious amounts of compute to train [56] or limit themselves to relatively small models [81, 37]. Contrary, MIM models have shown success and scalability to large models with comparatively little compute [29, 6, 64]. MIM-Refiner can build on these models which allows us to scale to large model sizes without a large amount of compute. Our largest model, MAE-Refined-2B contains approximately twice the parameters of the currently largest uni-modal contrastive model DINOv2-g [56] and can be trained on two orders of magnitude less data.
Attempts to preserve the efficiency while training MIM and ID objectives end-to-end have been less successful [47, 40], where both works came to the conclusion that sequential training (MIM $\rightarrow$ ID) circumvents various problem with end-to-end training. First, a powerful encoder is trained solely with a MIM objective. Second, the encoder is trained with a ID objective while preserving the rich features in early blocks with a layer-wise learning rate decay [17] in lower blocks and either constraining changes in early blocks [40] or completely freezing them [47].
MIM-Refiner is also a sequential MIM $\rightarrow$ ID method. In contrast to our work, previous works start from the representation after the last MIM encoder block and are therefore highly reliant on a good representation thereof. This can be seen for example on MAE-CT [47] where their ViT-H/14 model is worse than their ViT-H/16 model, despite using 30% more FLOPS. Additionally, previous works [47, 40] omit current go-to techniques such as multi-crop augmentation [10] which has been shown to improve performance [11, 82, 81].
Training models with additional losses from intermediate layers dates back to the early deep learning days [46, 65] where these auxiliary losses were used to alleviate optimization issues in lower layers. MIM-Refiner relates to deep supervision in the sense that we use multiple ID heads attached at intermediate layers where each head produces a loss for training.
5 Experiments
We refine a series of MIM models, namely MAE [29, 64], data2vec 2.0 [6] (abbreviated as D2V2), dBOT [50] and CrossMAE [27]. These models were pre-trained on ImageNet-1K, which we also use for the refinement. For visual clarity, we compare our best model against previous state-of-the-art models and show the full result tables in the Appendix. We evaluate our models in classification (low- and many-shot), feature evaluation, clustering, transfer learning and semantic segmentation.
5.1 Ablation Study
We first investigate various design choices of our method in Table 4(f).
(a) Adding multiple heads at intermediate features improves $k$-NN accuracy by 0.8%. The best settings include the blocks with the best representation quality (see Figure 3(b)). Adding additional heads before the best blocks does not improve performance while increasing training costs. (d) The benefit doubles to 1.6% for a ViT-H since deeper models degrade more (see Figure 3(b)). Additionally, we investigate where to attach the ID heads in Appendix C.2 where we find that simply attaching ID head to the last 8 (or 12 for ViT-H) blocks to perform best across model scales.
(b) Freezing early blocks as a form of regularization to preserve MIM features (similar to related works [47, 40]) is not necessary. Note that we still use a layer-wise learning rate decay [17].
(c) Scheduling the loss weight of each intermediate head is not necessary, i.e.  simply summing all losses is sufficient to achieve good performances. “Uniform Decay” decays the loss weight of all intermediate heads during training. “Staggered Decay” starts by decaying the first head and gradually starts to decay more and more heads as training progresses. “Staggered Step” disables the first head after some time followed by gradually disabling more and more heads. “One Hot” trains the encoder with only one intermediate head at a time where training starts with the earliest intermediate head and gradually iterates through the heads. Details to the loss schedules are in Appendix E.13.
(e) Using the NN swap only for aligning the positive with the anchor is slightly better than also swapping the negatives with their NN.
(f) Relying only on the data-driven augmentation of the NN-swap [24] by omitting color/blur augmentations, is beneficial for certain downstream tasks such as extreme low-shot classification [47]. We show more results without color/blur augmentations in Appendix C.3.
Table 1: Ablation study by refining data2vec 2.0 ViT-L/16 unless noted otherwise. Default settings
#Heads $k$-NN
1 80.2
2 80.2
4 80.9
8 81.0
12 81.0
#Frozen $k$-NN
0 81.0
6 80.9
12 80.6
Schedule $k$-NN
Constant 81.0
Uniform Decay 80.9
Staggered Decay 81.0
Staggered Step 80.9
One Hot 80.7
#Heads $k$-NN
1 80.7
12 82.3
Swap Neg. $k$-NN
81.0
80.9
Color/blur $k$-NN 1-shot
81.0 61.7
80.5 63.4
5.2 Low-shot and Feature Evaluations
We evaluate the ability of our models to perform low-shot classification in Table 2. Additionally, linear probing and $k$-NN accuracy are reported which are computed from the features of the frozen encoder. These metrics are typically correlated to low-shot performance as it indicates that the representation is already linear separable which eases drawing decision boundaries given only few labels. For linear probing, we use the protocol of DINOv2 [56] which includes using features of the last four blocks in its hyperparameter grid. Therefore, linear probing evaluates the feature representation at the end of the encoder, while $k$-NN accuracy evaluates only the features of the last block.
MIM-Refiner drastically improves upon MIM models and other SSL models. In the 1-shot settings D2V2-Refined-H sets a new state-of-the-art of 64.2%, outperforming the 63.6% of MAWS-6.5B [64] which is pre-trained on a private dataset with approximately 2000 times the size of ImageNet-1K.
Table 2: Low-shot and feature evaluations of recent SSL models on ImageNet-1K. MIM-Refiner significantly improves upon MIM models and previous state-of-the-art SSL models. Appendix Table 8 extends this comparison to more methods and to models that were trained on more data (such as DINOv2) where MIM-Refiner outperforms DINOv2-g in some benchmarks.
Low-shot Evaluation Feature Eval
ViT Method 1-shot 2-shot 5-shot 1% 10% Probe $k$-NN
L/16 MAE [29] 14.3 34.9 56.9 67.7 79.3 77.5 60.6
D2V2 [6] 24.1 58.8 72.1 75.1 81.5 78.2 51.8
iBOT [81] 48.5 58.2 66.5 73.3 79.0 81.1 78.0
Mugs [82] 52.9 62.3 69.4 76.2 80.3 82.1 80.4
MAE-Refined 57.8 66.3 72.0 76.1 81.2 82.8 81.5
D2V2-Refined 61.7 69.6 73.9 78.1 82.1 83.5 81.0
H/14 MAE [29] 7.2 14.1 40.2 72.8 81.2 78.2 58.1
D2V2 [6] 21.6 60.8 74.2 77.6 83.3 80.4 48.0
MAE-CT [47] 49.4 59.6 67.4 74.4 81.7 82.3 79.1
MAE-Refined 59.5 68.5 73.8 77.4 82.1 83.7 82.3
D2V2-Refined 64.2 71.3 75.5 78.1 83.5 84.7 82.3
2B/14 MAE [29, 64] 17.8 29.1 62.9 73.6 82.0 79.7 67.1
MAE-Refined 58.2 68.6 74.8 78.7 82.5 84.5 83.2
5.3 Cluster Evaluations
Table 3: $k$-means cluster performance and class separation on ImageNet of recent SSL models. MIM-Refiner drastically improves performance of MIM-models and even outperforms DINOv2-g which has 2x more parameters and is trained on on 100x more data.
Cluster Performance Class Separation
ViT Method ACC NMI AMI ARI SIL ($\uparrow$) DBS ($\downarrow$)
L/16 D2V2 [6] 10.5 45.1 19.5 2.5 -9.1 6.4
iBOT [81] 52.2 80.5 67.0 33.4 13.3 3.5
Mugs [82] 54.3 78.6 65.5 22.4 14.9 3.3
D2V2-Refined 67.4 86.3 76.2 40.5 37.1 2.2
H/14 D2V2 [6] 9.9 45.8 18.0 2.6 -10.8 6.5
D2V2-Refined 67.3 87.2 77.9 42.2 34.5 2.3
H/16 MAE-CT 58.0 81.8 69.3 36.8 - -
g/14 DINOv2 [56] (LVD-142M) 47.7 76.3 63.6 5.1 30.4 2.8
We compare the cluster performance of MIM-Refiner against recent SSL models in Table 3. We apply mini-batch $k$-means [63] 100 times to the validation set of ImageNet-1K and select the run with the lowest $k$-means loss for comparison (average performance is reported in Appendix Table 10). We report commonly used metrics for measuring Cluster Performance w.r.t. the ground truth clustering: Cluster Accuracy (ACC) [77], Normalized Mutual Information (NMI) [45], Adjusted Mutual Information (AMI) [54], Adjusted Rand Index (ARI) [38], where higher values indicate a better match of the found clustering with the ground truth labels. Further, we measure the Class Separation in the embedded space using the Davies-Bouldin score (DBS) [19] and silhouette score (SIL) [62] w.r.t. the ground truth classes. The DBS measures the separation and compactness of classes, where lower values are better and zero is the minimum. The SIL ranges from -100 to 100, where higher values are better. Negative SILs relate to mismatches between classes and embedding, where scores close to zero indicate a potential overlap of classes.
Table 3 shows that MIM-Refiner greatly improves various clustering metrics. The reached ACC of 67.4% outperforms the current state-of-the-art of 61.6% reached by TEMI MSN [1]. Figure 5 illustrate this drastic increase in cluster performance and class separation visually.
[…]
(a) D2V2
[…]
(b) D2V2-Refined
[…]
(c) Mugs
Figure 5: UMAP [53] plots of ViT-L embeddings using all 53 food related classes of ImageNet. The corresponding $k$-means cluster accuracy (ACC) and class separation measured in silhouette score (SIL) is shown below each plot. The clustering after refinement (b) is visually more condensed and better separated with corresponding improvements in ACC and SIL than before refinement (a). Mugs (c) does not separate the clusters that well, as shown by the merged clusters in the middle and the lower SIL score. The colors show the 53 ground truth food classes.
5.4 Transfer Learning Evaluations
We investigate generalization of pre-trained MIM-Refiner models to other datasets in Table 4, which shows the benefits of MIM-Refiner models when transferring a pre-trained representation. We consider a variety of classification downstream tasks and a semantic segmentation task. MAE-Refined consistently improves over MAE and state-of-the-art SSL methods. Additionally MIM-Refiner further improves when scaling up to a 2B parameter model.
Table 4: Transferring MIM-Refiner models to other datasets. “VTAB-6” reports the average accuracy over six datasets from the VTAB benchmark [78] and “VTAB-1K” is the average over all 19 datasets of the VTAB-1K benchmark. ADE20K reports the mean intersection over union (mIoU) of a semantic segmentation probe. MIM-Refiner learns general features that can easily be transferred to various datasets and tasks. Table 11 confirms this finding on additional models and Appendix C.7 shows individual VTAB performances for VTAB-6 probing and VTAB-1K fine-tuning.
iNat18 Fine-tuning Linear Probe VTAB-1K
ViT Method 1-shot 5-shot 10-shot iNat18 VTAB-6 ADE20K Fine-tuning
L/16 MAE 7.1 51.6 68.9 42.8 83.0 33.6 73.7
iBOT 15.8 51.5 65.5 56.0 87.6 35.6 70.8
Mugs 19.5 53.2 66.9 61.5 87.9 34.8 68.0
MAE-Ref. 19.0 58.0 71.7 60.6 88.5 37.3 75.2
H/14 MAE 6.5 53.3 71.7 43.0 83.2 35.5 72.7
MAE-CT 16.5 60.1 74.7 62.8 88.8 37.6 75.7
MAE-Ref. 20.9 62.4 75.4 64.6 89.3 39.4 75.9
2B/14 MAE 10.0 53.7 72.2 51.0 85.4 37.3 74.1
MAE-Ref. 22.5 63.5 76.5 69.6 89.8 40.3 75.6
5.5 Fine-tuning with Large Amounts of Labels
MIM models typically outperform ID methods when given enough labels to fine-tune the model. As our refinement process employs an ID objective, we investigate whether MIM-Refiner involuntarily degrades performance given an abundance of labels. To this end, we fine-tune MIM models and their refined version on ImageNet-1K [20], iNat18 [35] and ADE20K [80] using a linear classification/segmentation head. Table 5 shows a small but consistent improvent of MIM-Refiner models.
Table 5: Full fine-tuning using 100% of the labels. MIM-Refiner consistently improves performance slightly even though ID methods typically perform worse than MIM models on this benchmark. Appendix Table 14 confirms this pattern on additional MIM models. “IN-Rob.” shows the average performance of the trained IN-1K classifier on robustness datasets (individual results in Table 15).
ViT-L/16 ViT-H/14
Model IN-1K IN-Rob. iNat18 ADE20K IN-1K IN-Rob. iNat18
MAE [29] 85.7 53.9 80.7 47.6 86.7 60.3 82.7
MAE-Ref. 85.6 54.4 80.9 48.3 86.9 60.9 83.3
dBOT [50] 85.8 55.3 81.9 51.0 87.1 63.7 84.1
dBOT-Ref. 85.9 55.3 82.1 51.1 87.1 64.0 84.5
Mugs [82] 85.2 46.4 76.9 48.5 - -
6 Conclusion
We introduce MIM-Refiner, a procedure to refine pre-trained MIM models. Motivated by the insights that the representation quality of MIM models peaks in the middle of the encoder, we employ an ensemble of instance discrimination heads attached at multiple blocks towards the end of the encoder, including the blocks where representation quality peaks, to improve upon the best existing representation. We train this ensemble for a short duration to improve the representation for downstream tasks such as classification, clustering or semantic segmentation. As instance discrimination objective we introduce Nearest Neighbor Alignment, an adaption of existing nearest neighbor based objectives that uses the nearest neighbor relation only for aligning positive pairs and not for pushing negative pairs apart, to improve the signal quality from negatives.
Our refined MIM models learn strong features from ImageNet-1K alone that can be readily used for downstream tasks without fine-tuning the model but also improve performance when the model is fine-tuned, particularly in few-shot settings. Our models outperform competitors that were also trained on ImageNet-1K and sometimes also ones that were trained on more data or use larger models.
7 Acknowledgements
We thank Johannes Lehner for helpful discussions and suggestions.
We acknowledge EuroHPC Joint Undertaking for awarding us access to Karolina at IT4Innovations, Czech Republic, Leonardo at CINECA, Italy, MeluXina at LuxProvide, Luxembourg and LUMI at CSC, Finland.
The ELLIS Unit Linz, the LIT AI Lab, the Institute for Machine Learning, are supported by the Federal State Upper Austria. We thank the projects Medical Cognitive Computing Center (MC3), INCONTROL-RL (FFG-881064), PRIMAL (FFG-873979), S3AI (FFG-872172), EPILEPSIA (FFG-892171), AIRI FG 9-N (FWF-36284, FWF-36235), AI4GreenHeatingGrids (FFG- 899943), INTEGRATE (FFG-892418), ELISE (H2020-ICT-2019-3 ID: 951847), Stars4Waters (HORIZON-CL6-2021-CLIMATE-01-01). We thank Audi.JKU Deep Learning Center, TGW LOGISTICS GROUP GMBH, Silicon Austria Labs (SAL), FILL Gesellschaft mbH, Anyline GmbH, Google, ZF Friedrichshafen AG, Robert Bosch GmbH, UCB Biopharma SRL, Merck Healthcare KGaA, Verbund AG, GLS (Univ. Waterloo), Software Competence Center Hagenberg GmbH, Borealis AG, TÜV Austria, Frauscher Sensonic, TRUMPF and the NVIDIA Corporation.
Appendix A Reproducibility
We publish the full codebase used for this work which contains code for training the models, evaluations and dataset preparation. https://github.com/ml-jku/MIM-Refiner
Appendix B Limitations
One limitation of MIM-Refiner is that it requires batch normalization [39] layers in the ID head. Without them, performance decreases significantly. As we do not change anything in the MIM pre-training, we hypothesize that feature statistics after MIM pre-training are not suited for an ID task and therefore require per-feature normalization. The batch normalization layers significantly decrease scalability to distributed hardware setups as each layer requires a synchronization of batch statistics.
MIM-Refiner addresses a common issue with MIM models: their typically lightweight decoder often delegates part of the reconstruction to the encoder, resulting in subpar representations for the later encoder blocks in downstream tasks. Alternatively, one could simply argue for a larger decoder. However, a larger decoder increases computational costs since the decoder typically operates on the full sequence length. Additionally, the direction of a larger decoder was explored to a certain extent in the original MAE paper [29], where larger decoders performed worse in fine-tuning and linear probing. Successor models such as CrossMAE perform well with a deeper decoder but still show decreasing representation quality in later layers, as shown in Appendix C.1.
During early development, we tried various ways to propagate the intermediate representation towards the end. While we found that a simple ensemble of contrastive heads attached to later blocks works very well, there might be even better ways to leverage the rich intermediate MIM representations. We explored the following variants in with little success: (i) completely deleting the last few blocks (ii) reinitializing the last few blocks while setting the weights/biases of the last projection in the attention/MLP to 0 which leads to the result of the previous block being propagated to the end via the residual connection (iii) gating the last few blocks via ReZero [5].
Appendix C Extended Results
C.1 Intermediate Representation Analysis of Additional MIM Methods
We visualize the $k$-NN accuracies of various MIM models in Figure 6 where all of them show the pattern that the representation quality of larger models degrades towards the end of the encoder.
(a) MAE [29]
(b) data2vec 2.0 [6]
(c) CrossMAE [27]
(d) dBOT [50]
(e) CAE [15]
Figure 6: Intermediate representation analysis of various MIM models. As models get bigger, the $k$-NN accuracy degrades more towards the end of the encoder, especially for L/16 and H/14 models. This pattern is consistent over various MIM models.
C.2 Where to Attach ID Heads?
Table 6(b) ablates different choices of where to attach ID heads on a D2V2 pre-trained ViT-L/16 and ViT-H/14. On ViT-L/16, there is almost no difference of where to attach the ID heads in the last third of blocks. We tried to transfer this insight to ViT-H/14 where the default setting of attaching ID heads to all later blocks performs better. We therefore use the default setting of attaching ID heads to the last 8 blocks (ViT-L and ViT-2B) or the last 12 blocks (ViT-H).
Table 6: Spacing heads out more across the later ViT blocks can achieve comparable performances for ViT-L/16 but does not generalize to ViT-H/14. The default setting of attaching ID heads to all later blocks generalizes well across model scales.
Block Indices $k$-NN
20,24 80.9
16,20,24 81.1
15,18,21,24 81.1
18,20,22,24 81.0
17-24 81.0
(a) ViT-L/16
Block Indices $k$-NN
22,24,26,28,30,32 82.0
16,18,20,22,24,26,28,30,32 82.1
20-32 82.3
(b) ViT-H/14
C.3 Low-shot Classification Without Color Augmentations
MAE-CT [47] showed that omitting color augmentations can lead to performance gains for low-shot classification, especially on larger models. We therefore train our ViT-H/14 models also without color augmentations. We use the same hyperparameters as with color augmentations (see Table 18) except that we disable color augmentations (i.e. we train with only crop & flip augmentations) and half the training duration. Table 7 confirms the findings of [47] as omitting color augmentations also improves low-shot performance of MIM-Refiner.
Table 7: ImageNet-1K low-shot and feature evaluations of a refined data2vec 2.0 ViT-H/14 with and without color augmentations. Omitting color augmentations (i.e. using only crop & flip augmentations) improves ImageNet-1K low-shot performance.
Low-shot Evaluation Feature Eval
Model Color/blur 1-shot 2-shot 5-shot 1% 10% Probe $k$-NN
D2V2-Refined 64.7 72.0 75.9 79.1 83.5 84.1 82.1
64.2 71.3 75.5 78.1 83.5 84.7 82.3
C.4 Extended ImageNet-1K Low-shot and Feature Evaluations
Table 8 extends Table 2 with additonal results and comparison to more SSL models.
Table 8: Extending Table 2 with additional MIM-Refiner models and more SSL models. The last row-group compares the best MIM-Refiner model to models with longer sequence length (MSN-L/7) and models that were pre-trained on more data. Parentheses show pre-training dataset size. MIM-Refiner consistently improves MIM models, outperforms other SSL models with the same pre-training data and even outperforms DINOv2-g/14 in some settings.
Low-shot Evaluation Feature Eval
ViT Method 1-shot 2-shot 5-shot 1% 10% Probe $k$-NN
L/16 CrossMAE [27] 16.8 34.0 52.4 63.2 77.7 74.4 53.4
MAE [29] 14.3 34.9 56.9 67.7 79.3 77.5 60.6
dBOT [50] 28.0 46.9 62.4 70.0 80.2 77.8 61.3
D2V2 [6] 24.1 58.8 72.1 75.1 81.5 78.2 51.8
CAE [15] 28.1 56.5 68.4 71.4 79.5 80.0 66.9
iBOT [81] 48.5 58.2 66.5 73.3 79.0 81.1 78.0
Mugs [82] 52.9 62.3 69.4 76.2 80.3 82.1 80.4
CrossMAE-Refined 50.3 60.9 68.2 71.7 79.3 81.8 79.9
MAE-Refined 57.8 66.3 72.0 76.1 81.2 82.8 81.5
dBOT-Refined 57.4 66.0 71.7 76.6 81.6 83.3 81.3
D2V2-Refined 61.7 69.6 73.9 78.1 82.1 83.5 81.0
H/14 MAE [29] 7.2 14.1 40.2 72.8 81.2 78.2 58.1
dBOT [50] 23.9 45.0 63.0 73.0 82.1 79.0 60.0
D2V2 [6] 21.6 60.8 74.2 77.6 83.3 80.4 48.0
I-JEPA [3] 35.1 47.9 59.9 73.3 79.5 79.3 71.6
MAE-CT [47] 49.4 59.6 67.4 74.4 81.7 82.3 79.1
MAE-Refined 59.5 68.5 73.8 77.4 82.1 83.7 82.3
dBOT-Refined 59.2 67.6 72.9 77.3 82.5 84.0 82.0
D2V2-Refined 64.2 71.3 75.5 78.1 83.5 84.7 82.3
2B/14 MAE [29, 64] 17.8 29.1 62.9 73.6 82.0 79.7 67.1
MAE-Refined 58.2 68.6 74.8 78.7 82.5 84.5 83.2
L/16 iBOT [81] (14M) 37.4 49.9 61.9 70.9 80.3 82.7 72.9
L/7 MSN [2] (1.3M) 57.1 66.4 72.1 75.1 - 80.7 -
H/14 D2V2-Refined (1.3M) 64.2 71.3 75.5 78.1 83.5 84.7 82.3
g/14 DINOv2 [56] (142M) 60.5 68.3 74.4 79.1 83.8 86.5 83.5
C.5 Extended ImageNet-1K Cluster Evaluations
Table 9 extends the main paper results (Table 3) with additional models. Table 10 shows the average clustering results of the 100 mini-batch $k$-means runs as described in Section 5.3. We see that MIM-Refiner has also better average performance than competitor methods.
Table 9: Best $k$-means cluster performance and class separation on ImageNet of recent SSL models. This table extends Table 3 from the main paper with additional comparison and refined models.
Cluster Performance Class Separation
ViT Method ACC NMI AMI ARI SIL ($\uparrow$) DBS ($\downarrow$)
L/16 MAE [29] 18.5 55.2 29.1 6.9 -5.9 5.0
D2V2 [6] 10.5 45.1 19.5 2.5 -9.1 6.4
iBOT [81] 52.2 80.5 67.0 33.4 13.3 3.5
Mugs [82] 54.3 78.6 65.5 22.4 14.9 3.3
MAE-Refined 61.8 84.0 72.6 40.7 21.4 2.9
D2V2-Refined 67.4 86.3 76.2 40.5 37.1 2.2
H/14 MAE [29] 14.3 50.2 24.2 4.3 -7.8 5.2
D2V2 [6] 9.9 45.8 18.0 2.6 -10.8 6.5
MAE-Refined 64.6 85.3 74.6 45.5 21.0 2.9
D2V2-Refined 67.3 87.2 77.9 42.2 34.5 2.3
2B/14 MAE [29] 19.9 54.1 33.1 6.2 -3.6 4.8
MAE-Refined 63.0 85.0 74.4 44.0 14.0 3.2
g/14 DINOv2 [56] 47.7 76.3 63.6 5.1 30.4 2.8
Table 10: Average $k$-means cluster performance on ImageNet of recent SSL models. MIM-Refiner drastically improves performance of unrefined models and outperforms competitors.
Cluster Performance
ViT Method ACC NMI AMI ARI
L/16 MAE [29] 17.9 54.5 28.9 6.5
D2V2 [6] 10.2 44.5 19.3 2.3
iBOT [81] 50.5 80.0 66.6 31.6
Mugs [82] 50.7 77.4 64.4 18.1
MAE-Refined 60.6 83.5 71.9 35.0
D2V2-Refined 60.6 83.9 72.9 30.0
H/14 MAE [29] 13.8 49.8 24.4 4.2
D2V2 [6] 9.7 45.2 18.1 2.5
MAE-Refined 63.2 84.7 74.0 40.1
D2V2-Refined 60.4 84.5 74.3 28.4
2B/14 MAE [29, 64] 19.2 53.5 32.9 5.7
MAE-Refined 59.4 84.0 73.4 38.0
g/14 DINOv2 [56] 46.8 75.6 62.7 3.5
Figure 7 shows an additional analysis on the cluster structure in each of the blocks for the refined and unrefined MAE and D2V2 models with ViT-H/14 on ImageNet-1K. Corresponding to the $k$-NN accuracy analysis in Figure 3(b), we see that MAE and D2V2 have a higher cluster accuracy (ACC) in the intermediate blocks than in the last block. MIM-Refiner turns this behaviour around and causes a steep increase of ACC starting from the intermediate block and continuing to almost 70% in the last layer. The silhouette score (SIL) confirms this as well. The early blocks allow no separation of the ground truth ImageNet-1K classes leading to negative SIL values, whereas later blocks of the refined models increase separation by 30-50% compared to unrefined counterparts. Interestingly, MAE-Refined is not plateauing in SIL and ACC compared to D2V2-Refined, pointing to potential room for improvement in MAE refinement by using additional ID heads at earlier layers. The bottom part of Figure 7 measures the pairwise cluster label similarity between subsequent blocks in terms of normalized mutual information (NMI) [45] as $\left(\operatorname{NMI}(y_{i},y_{i+1})\right)\cdot 100$, where $y_{i}$ are the $k$-means cluster labels at block $i$. The low cluster label similarity in the early blocks for all models indicates that the found cluster labels focus on different clusterings. The later blocks of MIM-Refined models have high alignment with similarities of more than 90%. The higher ACC w.r.t. the ground truth ImageNet-1K classes in the later blocks indicates that they focus more on object-centric features. This is in contrast to the overall lower cluster label similarity in unrefined models, which focus on different cluster structures in each block.
Figure 7: Block-wise cluster analysis for refined and unrefined MAE and D2V2 (H/14). Upper Left: Cluster accuracy w.r.t. lowest $k$-means loss per block. Upper Right: Silhouette score w.r.t. ground truth ImageNet-1K classes. Bottom: Pairwise cluster label similarity between subsequent blocks (details in Section C.5).
C.6 Extended Transfer Learning Results
Table 11 extends Table 4 with additional results and comparison to more SSL models.
Table 11: Extending Table 4 with additional MIM-Refiner models and additional SSL models. MIM-Refiner learns general features that can easily be transferred to various datasets and tasks.
iNat18 Fine-tuning Linear Probe
ViT Method 1-shot 5-shot 10-shot iNat18 VTAB-6 ADE20K
L/16 CrossMAE [27] 5.4 45.6 63.7 39.8 81.2 30.9
MAE [29] 7.1 51.6 68.9 42.8 83.0 33.6
dBOT [50] 7.1 53.1 71.2 44.6 83.4 34.1
D2V2 [6] 5.5 53.1 70.6 39.4 81.5 38.8
CAE [15] 7.4 54.5 71.2 46.4 84.8 36.5
iBOT [81] 15.8 51.5 65.5 56.0 87.6 35.6
Mugs [82] 19.5 53.2 66.9 61.5 87.9 34.8
CrossMAE-Ref 18.0 52.9 67.8 60.5 88.3 34.3
MAE-Ref 19.0 58.0 71.7 60.6 88.5 37.3
dBOT-Ref 18.3 58.8 73.0 61.7 88.5 38.4
D2V2-Refined 15.2 56.3 71.8 52.0 85.9 41.0
H/14 MAE [29] 6.5 53.3 71.7 43.0 83.2 35.5
dBOT [50] 6.8 53.1 73.2 46.2 84.6 36.1
D2V2 [6] 5.9 55.7 73.4 41.7 83.1 42.4
MAE-CT [47] 16.5 60.1 74.7 62.8 88.8 37.6
MAE-Ref 20.9 62.4 75.4 64.6 89.3 39.4
dBOT-Ref 20.0 60.9 75.8 65.8 89.5 37.6
D2V2-Refined 16.1 59.2 74.8 54.4 87.1 43.7
2B/14 MAE [64] 10.0 53.7 72.2 51.0 85.4 37.3
MAE-Ref 22.5 63.5 76.5 69.6 89.8 40.3
L/7 MSN [2] 17.0 38.0 48.1 - - -
C.7 VTAB Individual Dataset Results
We show the individual accuracies for each VTAB [78] dataset from Table 4 of the main paper. Table 12 shows linear probing results on VTAB and Table 13 shows fine-tuning results on VTAB-1K. We use only six datasets for linear probing, as the other datasets are quite different to the ImageNet-1K images seen during training and therefore benefit heavily from fine-tuning which makes them more suited for evaluation via fine-tuning instead of linear probing.
Table 12: Linear probing accuracy on six VTAB [78] datasets from the “Natural” category: Caltech101 [26], CIFAR100 [44], DTD [16], Flowers102 [55], Pets [57] and Sun397 [74].
VTAB Dataset
ViT Method CF100 CT101 DTD FL102 Pets Sun397 Average
L/16 CrossMAE 81.4 88.8 75.6 84.2 84.1 72.7 81.2
MAE 80.0 91.8 75.6 86.3 89.6 74.7 83.0
dBOT 84.3 91.3 75.7 88.0 86.0 75.1 83.4
D2V2 85.0 89.8 74.0 84.9 80.7 74.6 81.5
CAE 87.0 92.3 76.1 88.4 89.2 75.9 84.8
iBOT 89.3 91.3 78.1 95.8 93.7 77.1 87.6
Mugs 89.5 90.5 78.0 96.8 95.3 77.2 87.9
CrossMAE-Ref 88.7 93.5 79.0 95.7 95.2 78.1 88.3
MAE-Ref 89.1 91.9 79.0 96.2 95.8 78.8 88.5
dBOT-Ref 90.4 91.3 78.6 95.9 95.5 79.2 88.5
D2V2-Ref 88.9 88.8 73.9 92.1 94.7 77.1 85.9
H/14 MAE 81.0 90.3 76.9 85.9 89.5 75.3 83.2
dBOT 85.5 91.7 77.7 88.1 88.2 76.3 84.6
D2V2 87.1 91.4 77.4 85.9 79.9 76.9 83.1
I-JEPA 87.1 92.7 72.5 90.4 92.4 74.9 85.0
MAE-CT 87.7 93.9 80.1 97.0 95.0 79.2 88.8
MAE-Ref 90.1 92.0 80.4 97.5 96.0 79.8 89.3
dBOT-Ref 91.7 92.1 80.6 96.7 96.1 80.1 89.5
D2V2-Ref 90.4 89.0 75.9 93.3 95.6 78.4 87.1
2B/14 MAE 82.5 92.0 78.2 90.5 91.8 77.1 85.4
MAE-Ref 90.8 92.6 81.1 97.7 96.5 80.3 89.8
Table 13: Fine-tuning accuracy of all 19 VTAB-1K [78] datasets (averages are reported in Table 4). Row-groups correspond to ViT-L/16, ViT-H/14 and ViT-2B/14 models respectively.
Natural Specialized Structured
Method
CF100
CT101
DTD
FL102
Pets
SVHN
Sun397
Camelyon
EuroSAT
Resisc45
Retinopathy
Clevr-Count
Clevr-Dist
DMLab
KITTI-Dist
dSpri-Loc
dSpr-Ori
sNORB-Azi
sNORB-Ele
MAE 50.3 90.1 68.0 90.8 92.7 92.2 35.1 85.5 93.2 81.7 72.7 93.2 63.9 55.8 84.5 93.9 58.6 46.1 51.9
D2V2 55.6 91.8 66.8 92.5 91.4 88.6 38.1 85.5 92.7 82.9 73.6 95.7 64.5 52.4 84.7 95.6 58.9 45.8 57.1
iBOT 68.3 91.5 71.5 95.8 91.8 80.1 47.6 85.4 93.4 86.6 74.5 79.9 61.7 52.3 80.2 74.5 47.2 27.5 35.4
Mugs 69.4 89.2 71.4 95.4 92.9 78.1 47.6 85.2 94.0 86.2 74.7 75.1 53.0 51.1 77.9 46.8 45.5 26.3 32.6
MAE-Ref 62.2 90.6 71.5 97.2 93.7 91.7 49.1 85.3 94.8 86.8 73.4 92.5 64.3 57.4 84.3 88.4 60.5 39.0 46.5
D2V2-Ref 65.0 90.0 69.0 93.9 91.8 85.8 47.8 71.4 94.1 83.4 71.9 94.0 65.0 56.5 85.0 89.1 59.9 40.8 46.7
MAE 44.3 88.3 67.8 90.8 90.8 91.1 30.4 85.9 94.0 81.3 73.9 92.7 63.6 54.0 84.9 92.2 62.8 37.1 55.2
D2V2 61.0 92.0 72.3 94.7 92.6 90.2 41.1 84.9 93.3 85.3 73.6 97.3 64.4 54.3 84.1 96.6 57.6 42.0 55.2
MAE-CT 58.5 93.2 72.3 97.5 93.4 90.7 48.3 85.9 94.4 86.1 75.0 93.0 64.9 58.3 83.6 91.9 63.8 38.0 48.7
MAE-Ref 62.3 91.3 73.1 97.9 93.9 89.5 49.7 86.8 94.9 87.7 75.1 92.9 63.3 57.5 85.9 93.5 62.2 36.0 49.3
D2V2-Ref 64.9 89.0 66.5 94.1 92.3 88.9 45.9 84.9 62.6 82.3 73.6 95.3 56.6 57.7 84.6 92.5 59.7 36.9 38.5
MAE 44.7 90.4 70.0 91.8 92.4 92.3 33.8 85.9 94.2 82.6 74.2 92.4 63.9 55.5 85.5 94.6 61.1 44.9 57.9
MAE-Ref 61.8 91.0 73.9 97.5 94.5 91.3 49.1 86.6 94.7 87.0 73.8 92.6 64.7 58.8 85.3 93.5 61.2 37.9 40.9
C.8 Extended Fine-tuning Results
Table 14 extends Table 5 with additional MIM models. MIM-Refiner consistently improves also on fine-tuning with an abundance of labels. Table 15 shows individual performances on robustness datasets for fine-tuned models on ImageNet-1K.
Table 14: Full fine-tuning using 100% of the labels. MIM-Refiner consistently improves performance slightly even though ID methods typically perform worse than MIM models on this benchmark. This table extends Table 5 from the main paper with additional MIM models and SSL models.
ViT-L/16 ViT-H/14
Model ImageNet-1K iNat18 ImageNet-1K iNat18
MAE 85.7 80.7 86.7 82.7
MAE-Refined 85.6 80.9 86.9 83.3
CrossMAE 84.9 77.7 - -
CrossMAE-Refined 85.1 78.4 - -
D2V2 86.6 81.0 86.6 79.6
D2V2-Refined 86.7 81.6 86.8 79.8
dBOT 85.8 81.9 87.1 84.1
dBOT-Refined 85.9 82.1 87.1 84.5
iBOT 84.8 76.9 - -
Mugs 85.2 76.9 - -
I-JEPA - - 84.9 75.9
Table 15: Robustness and domain generalization evaluation on ImageNet-C(orruption) [32] ImageNet-A(dversarial) [34], ImageNet-R(endition) [31] and ImageNet-Sketch [71]. For ImageNet-C we report the mean corruption error [32]. MIM-Refiner consistently improves robustness, particularly on larger models. Table 5 reports the average accuracy of IN-A, IN-R and IN-Sketch.
Robustness Dataset
ViT Method IN-C ($\downarrow$) IN-A ($\uparrow$) IN-R ($\uparrow$) Sketch ($\uparrow$) Validation ($\uparrow$)
L/16 CrossMAE 41.0 51.8 57.1 42.0 84.9
CrossMAE-Ref. 40.7 52.1 57.0 42.1 85.1
L/16 MAE 39.2 56.2 60.0 45.6 85.7
MAE-Ref. 39.0 57.0 60.1 45.9 85.6
L/16 D2V2 34.0 66.1 64.4 50.0 86.6
D2V2-Ref. 33.5 67.2 64.3 50.1 86.7
L/16 dBOT 36.1 60.0 60.5 45.5 85.8
dBOT-Ref. 36.1 60.3 60.4 45.3 85.9
L/16 iBOT 37.8 47.6 53.7 41.9 84.5
Mugs 39.9 47.8 52.0 39.3 84.6
H/14 MAE 34.6 67.8 64.1 48.9 86.7
MAE-Ref. 34.3 68.0 65.1 49.7 86.9
H/14 D2V2 30.7 72.9 65.7 51.0 86.6
D2V2-Ref. 30.2 74.1 66.1 52.1 86.8
H/14 dBOT 31.8 71.1 68.1 51.8 87.1
dBOT-Ref. 31.8 72.3 68.0 51.8 87.1
H/14 I-JEPA 37.4 51.7 57.4 43.7 81.7
2B/14 MAE 32.6 68.4 65.4 50.0 86.7
MAE-Ref. 32.2 68.9 66.2 50.5 86.8
C.9 Fine-tuning ViT-2B Models
Table 16 shows results for fine-tuning ViT-2B models.
Table 16: Fine-tuning ViT-2B models using 100% of the labels. As fine-tuning these models is expensive, we freeze the first 6 of the 24 blocks to save memory and compute.
ViT-2B/14
Model ImageNet-1K iNat18
MAE 86.7 82.2
MAE-Refined 86.9 83.2
C.10 Impact of Multi-Crop Augmentation
Multi-crop augmentation [10] has been shown to greatly improve the performance of ID methods [10, 11, 81, 82] and also improves MIM-Refiner significantly, where the $k$-NN accuracy of a D2V2-Refined L/16 drops by 2.6% when omitting multi-crop augmentations. When comparing to drops of other models, this is a relatively small drop. For example, the performance of DINO B/16 drops by 7.2% and iBOT B/16 drops by 5.6% when omitting multi-crop augmentations (see Table 10 in [81]).
C.11 High-dimensional $k$-NN
The linear probing protocol of DINOv2 [56] includes the possibility to concatenate features of the last 4 blocks as input to the linear probe. We find that this outperforms using only the features of the last block in most cases. As the best linear probe uses features from the last 4 blocks and the fact that the $k$-NN and linear probing metrics are typically correlated [56], we report the $k$-NN accuracy using only the features of the last block in Table 2 to use the last 4 blocks for one metric and only the last block for the other. In Table 17 we investigate using the concatenation of features from the last 4 blocks for a $k$-NN classifier. Most models benefit from using more features, especially MIM models.
Table 17: ImageNet-1K $k$-NN accuracy at 224x224 resolution of the [CLS] token of the last block or the concatenation of the [CLS] tokens of the last 4 blocks.
#Blocks
ViT Method 1 4 Delta
L/16 MAE 60.6 63.3 +2.7
D2V2 51.8 52.9 +1.1
iBOT 78.0 78.9 +0.9
Mugs 80.4 80.1 -0.3
MAE-Refined 81.5 81.5 0.0
D2V2-Refined 81.0 81.7 +0.7
H/14 MAE 58.1 61.4 +3.3
D2V2 48.0 52.2 +4.2
I-JEPA 71.6 72.3 +0.7
MAE-CT 79.1 78.6 -0.5
MAE-Refined 82.3 82.5 +0.2
D2V2-Refined 82.3 83.4 +1.1
g/14 DINOv2 83.0 83.9 +0.9
Appendix D Implementation Details
D.1 Evaluations
To avoid slight performance differences due to minor implementation details or version changes and facilitate a fair comparison between models, we run all evaluations on our own in accordance with the suggested hyperparameters of the respecitve methods.
D.2 Hardware
All models are pre-trained on multiple nodes of 4xA100-64GB GPUs where ViT-L uses 4 nodes (i.e. 16 GPUs), ViT-H 8 nodes of 4xA100 (i.e. 32 GPUs) and ViT-2B uses 16 nodes (i.e. 64 GPUs). For evaluations, we use a mix of 4xA100-64GB nodes, 8xA100-40GB nodes and various smaller nodes that vary in number of GPUs. We estimate the total number of A100 GPU-hours used for this project to be 40K hours. This estimate includes everything from initial exploration, method development, analysis and evaluations.
D.3 Vision Transformer
The architecture of our models follows the ones from the respective MIM model that is refined. That is a pre-norm architecture for MAE [29] and a post-norm architecture for data2vec 2.0 [6]. We attach the ID heads to the [CLS] tokens and also use the [CLS] token for evaluation.
We download the official checkpoints from the respective MIM works. Note that the official CrossMAE model is pre-trained for less epochs than all other MIM models. For dBOT we use the models where a teacher of the same size is used (i.e. dBOT-L used MAE-L as teacher and dBOT-H used MAE-H as teacher).
D.4 ID Head Architecture
We use a three layer MLP with hidden dimension 2048 as projector and a two layer MLP with hidden dimension 4096 as predictor. Each linear projection is followed by a GELU [33] and a batchnorm [39] layer. For the last linear projection in projector and predictor, no GELU is used.
D.5 Refinement Hyperparameters
Hyperparameters for the refinement stage are listed in Table 18. Following MAE-CT [47], we initialize all ID heads first by training them with a frozen encoder to ensure a good learning signal from the start of the refinement process. For this initialization, we use the same hyperparameters as in Table 18 except that we use 20 epochs for all models, a learning rate of 2e-4 and a top1-NN lookup. As we do not use a momentum encoder during training, we instead track an EMA of the encoder and use the EMA then for downstream tasks. As ViT-2B is very expensive to train, we freeze the first 6 blocks (for refinement and also for evaluation). As shown in Table 4(b) this slightly reduces performance but also reduces memory consumption and runtime.
Table 18: MIM-Refiner hyperparameters.
Parameter Value
Epochs 30 (MAE L/H), 20 (MAE 2B, data2vec 2.0)
Batch Size 1024 (L), 512 (H, 2B)
Optimizer AdamW [42, 51]
   Momentum $\beta_{1}=0.9,\beta_{2}=0.95$
Learning Rate Schedule Linear Warmup $\rightarrow$ Cosine Decay
   Warmup Epochs 4
   Learning Rate 4e-4
Encoder
   Layer-wise LR Decay [17] 0.65
   Freeze Blocks 0 (everything exept ViT-2B), 6 (ViT-2B)
   Weight Decay 0.05
   EMA [60] 0.9999
NNCLR Heads
   Weight Decay 1e-5
   Temperature 0.2 (L), 0.3 (H), 0.35 (2B)
   topk-NN $k$ 20
   NN-swap for Positives
   NN-swap for Negatives
Data Augmentation
   Color & Blur Settings see BYOL [28]
   Global Views 2
   Global View Resolution 224
   Global View Scale [0.25, 1.0]
   Local Views 10
   Local View Resolution 96 (L), 98 (H)
   Local View Scale [0.05, 0.25]
Appendix E Evaluation Details
E.1 GPU Hours Benchmark
For benchmarking GPU hours, we follow the setup from MAE-CT [47]: we conduct a comparison by implementing all methods in a single code-base and conducting short training runs on a single A100 40GB PCIe card. These runs are executed in mixed precision training mode and with the highest possible batchsize that is a power of 2. The runtime of these benchmark runs is then extrapolated to the reported number of epochs. Benchmarks are conducted in pytorch 2.1 with CUDA 12.1. FLOPS are measured with the fvcore library [^1] https://github.com/facebookresearch/fvcore. For the 1-shot classification plot in Figure 2, we do not take into account that some models train on higher resolutions (e.g. DINOv2) for visual clarity.
E.2 MAE Intermediate Representation Analysis
We analyze how well the features of a ViT block are suited for reconstruction by training an MAE with a decoder attached after every ViT block. We use the same parameters as for training from scratch [29] but reduce training duration to 20 epochs, warmup to 5 epochs and the depth of all decoders to 2. The encoder remains fully frozen during training and only the decoders are trained.
For the visualization in Figure 3(d), we calculate the delta from one block to the next and normalize the deltas by dividing by the maximum delta. We do this for both the $k$-NN accuracy and the reconstruction loss. Figure 8 shows the reconstruction loss per block and the same plot for a MAE L/16, where a similar behavior can be observed.
Figure 8: Left: Reconstruction loss per block of MAEs. Right Relative improvement of reconstruction loss and $k$-NN accuracy for a MAE L/16.
E.3 ImageNet-1K Low-shot Evaluation Details
For the 1, 2 and 5-shot benchmarks we train a logistic regression [11, 2] using the [CLS] token after the last encoder block with the cyanure [52] library. As MIM models benefit from fine-tuning in this setting [2], MAE and data2vec 2.0 are fine-tuned instead. We report the average of three dataset splits from MSN [2].
In the 1% and 10% low-shot benchmark, all models are fine-tuned with hyperparameters similar to those used in related works [47, 3, 9]. As the parameters vary between 1%/10% and also between model sizes, we refer to the codebase for the exact protocols.
For a fair comparison, we conduct the low-shot evaluations of DINOv2 at 224 resolution. We study the impact of the higher resolution where we observe minimal gains at the original resolution (518x518). Note that DINOv2 first trains at 224x224 followed by a short training at 518x518 resolution.
Table 19: ImageNet-1K low-shot evaluation of DINOv2 g/14 on higher resolutions.
resolution #patches FLOPS [G] 1-shot 2-shot 5-shot
224x224 256 291 60.5 68.3 74.4
518x518 1369 1553 61.1 68.8 74.8
E.4 ImageNet-1K $k$-NN Classification Details
For $k$-NN classification, we follow the protocol of DINO [73, 11]. We train a soft $k$-NN classifier weighted by cosine similarity with $k=10$. For MIM models, higher values for $k$ are beneficial, so we tune this parameter for MAE and data2vec 2.0.
E.5 ImageNet-1K Linear Probing Evaluation Details
For linear probing, we use the protocol from DINOv2 [56] for publicly released models and the values from the original papers otherwise. We train for 50 epochs using SGD with momentum 0.9. As data augmentation we use $\tt{RandomResizedCrop}$ and $\tt{HorizontalFlip}$. The DINOv2 protocol sweeps over the following hyperparameters by training multiple linear classifiers at once:
  • 13 learning rates ranging from 0.0001 to 0.5
  • Use the last block output or concatenate the output of the last 4 blocks
  • Use the [CLS] token or the concatenation of [CLS] and [AVG] token
As the linear probes trained on the concatenation of the last 4 blocks have more features and more parameters to discriminate between classes, they tend to be the best within the sweeped parameters. Note that we evaluate the representation of the last block in isolation via $k$-NN classification. We investigate $k$-NN classification with features from the last 4 blocks in Appendix C.11.
E.6 ImageNet-1K Cluster Evaluation Details
For each considered model in the clustering experiments in Section 5.3 we used the CLS token embeddings of the ImageNet validation set and preprocessed the embeddings using L2 normalization. For conducting mini-batch $k$-means and calculating the cluster related metrics we used the scikit-learn package [59], except for calculating the cluster accuracy where we used the implementation in ClustPy [48]. The UMAP plots in Figure 5 where generated by applying UMAP on top of the L2 normalized CLS token embeddings of the 53 food related classes of ImageNet for each model. We use the default UMAP parameters of umap-learn [53] for all plots (n_neighbors=15).
E.7 iNat18 Transfer Learning Evaluation Details
We report the accuracy on the validation set averaged over three seeds.
For 1-shot classification on iNat18, we use the linear probing protocol from DINOv2 [56]. We also attempted to fine-tune MIM models where some models fail to exceed random performance and therefore also use linear probing.
For 5-shot and 10-shot classification on iNat18, we fine-tune all models. The hyperparameters for fine-tuning (Table 20) are inspired by MAWS [64].
Table 20: Hyperparameters for fine-tuning on iNat18 low-shot classification.
Parameter Value
Epochs 50
Batch size 256
Optimizer AdamW [51, 42]
   Learning rate 1e-3
   Layer-wise lr decay [17] 0.75
   Weight decay 0.05
   Momentum $\beta_{1}=0.9,\beta_{2}=0.999$
Learning rate schedule linear warmup $\rightarrow$ cosine decay
   Warmup epochs 5
Label smoothing [66] 0.1
Data Augmentation
   Resize 256
    interpolation bicubic
   RandomResizedCrop 224
    scale [0.08, 1.0]
    interpolation bicubic
   RandomHorizontalFlip $p=0.5$
   Normalize ImageNet statistics
E.8 Transfer Learning Linear Probing
For transfering the pre-trained features to iNat18, six VTAB datasets and ADE20K (Table 4) we use the DINOv2 [56] linear probing protocol as described in Appendix E.5. For ADE20K and iNat18, we reduce the hyperparameter grid to fit into 40GB GPU memory.
E.9 ADE20K Semantic Segmentation Linear Probe
Large models (such as ViT-H or ViT-2B) are expensive to train on ADE20K. Therefore, we opt for a simple light-weight evaluation protocol to compare our models on a segmentation task:
  • We keep resolution at 224x224
  • We freeze the encoder
  • We train a linear classifier similar to DINOv2 [56] that predicts a class for each patch. The resulting low-resolution prediction is then upsampled to 224x224 resolution.
  • For evaluation, we use the original resolution image and slide a 224x224 window over the image with a stride of 170 pixels and average the logits per pixel.
As intermediate representations are commonly used for semantic segmentation, we use features from the last block, the 5th last block, the 9th last block and the 13th last block. Compared to simply using the last 4 blocks, this improves performance for all compared models.
E.10 Fine-tuning on VTAB-1K
For fine-tuning models on VTAB-1K we provide the hyperparameters in Table 21. We search for the best learning rate for each dataset by fine-tuning the model 25 times (5 learning rates with 5 seeds each) on the 800 training samples and evaluating them on the 200 validation samples. With the best learning rate, we then train each model 5 times on concatenation of training and validation split, evaluate on the test split and report the average accuracy.
Table 21: Hyperparameters for fine-tuning on VTAB-1K.
Parameter Value
Epochs 50
Batch size 64
Seeds 5
Optimizer AdamW [51, 42]
   Learning rate [1e-3, 7.5e-4, 5.0e-4, 2.5e-4, 1.0e-4]
   Layer-wise lr decay [17] 0.75
   Weight decay 0.05
   Momentum $\beta_{1}=0.9,\beta_{2}=0.999$
Learning rate schedule linear warmup $\rightarrow$ cosine decay
   Warmup epochs 5
Data Augmentation
   Resize
    interpolation bicubic
    size 224x224
   Normalize ImageNet-1K statistics
E.11 Fine-tuning With 100% Labels
For fine-tuning with 100% of the labels (Table 5), we use the hyperparameters provided in MAE [29] for both iNat18 and ImageNet-1K (see Table 22). As D2V2 models are unstable with the default learning rate of the MAE fine-tuning protocol, we use the highest stable learning rate out of 5e-4, 2.5e-4 and 1e-4.
For ViT-2B/14 (Table 16), we freeze the first 6 of the 24 blocks to reduce computational costs. Additionally, as the 2B models are sometimes unstable with a learning rate of 1e-3, we reduce it to 7.5e-4 or 5e-4 using the largest stable learning rate.
To fine-tune I-JEPA [3], we adjust hyperparameters to match their fine-tuning setting of a ViT-H/16448. We reduce peak stochastic depth from $0.3$ to $0.25$. To fine-tune on iNat18, we found that a learning rate 1e-3 performs better than the 1e-4 used in the original work.
Table 22: Hyperparameters for fine-tuning on ImageNet-1K and iNat18 many-shot classification.
Parameter Value
Epochs 50
Batch size 1024
Stochastic depth [36]
   Peak rate 0.2 (ViT-L/2B), 0.3 (ViT-H)
   Decay
Optimizer AdamW [51, 42]
   Learning rate 1e-3
   Layer-wise lr decay [17] 0.75
   Weight decay 0.05
   Momentum $\beta_{1}=0.9,\beta_{2}=0.999$
Learning rate schedule linear warmup $\rightarrow$ cosine decay
   Warmup epochs 5
Label smoothing [66] 0.1
Train Data Augmentation
   RandomResizedCrop 224
    scale [0.08, 1.0]
    interpolation bicubic
   RandomHorizontalFlip $p=0.5$
   RandAug [18]
    magnitude 9
    magnitude_std 0.5
   Normalize ImageNet statistics
   Mixup [79] $\alpha$ 0.8
   Cutmix [79] $\alpha$ 1.0
Test Data Augmentation
   Resize 256
    interpolation bicubic
   CenterCrop 224
   Normalize ImageNet statistics
E.12 ADE20K Semantic Segmentation Fine-tuning
We fine-tune ViT-L models using a linear head to predict a class label per 16x16 patch which is then upscaled to the original resolution. We follow common practices and fine-tune on 512x512 resolution, where we interpolate the absolute positional embedding from 224x224 to 512x512, add relative position bias to the attention layers (initialized to 0) [29] and introduce layerscale [67] (initialized to 1). A common augmentation pipeline is used that consists of random rescaling, random horizontal flipping, color jitter and padding if necessary. We train for 40K updates using a batchsize of 16, a learning rate of 1e-4, weight decay 0.05, linear warmup for 1.5K update steps followed by cosine decay, stochastic depth rate 0.2, layer-wise learning rate decay 0.95. We evaluate after 40K update steps using a sliding window of 341 pixels and report mIoU over the validation set.
This setting closely follows [41], who showed that the performance of this relatively simple pipeline has a high correlation to performance larger models and more sophisticated segmentation pipelines.
E.13 Loss Schedule Visualizatios
We visualize the schedules used for scheduling the loss weight of ID heads attached at intermediate blocks in the ablation study (Table 4(f)) in Figure 9.
Figure 9: Loss weight schedules for the ablation in Table 4(f). For visual clarity, small offsets are added when values overlap (for “One Hot” and “Staggered Step”).
Appendix F Relation to NNCLR
Figure 10 shows the difference between NNCLR [24] and NNA. NNCLR uses the NN-swap also for the negatives, resulting in a worse signal due to the NNs being retrieved from a FIFO queue of features from previous model states.
Figure 10: The NN-swap of NNCLR introduces inter-sample correlations between positives but uses features from a previous state of the model. Using the NN-swap only for the positives preserves the inter-sample correlations while using features from the current state of the model as negatives to improve the loss signal.
Appendix G Practitioner’s Guide
We find MIM-Refiner to be easy to tune. By freezing the encoder and training multiple ID heads attached to the encoder with different hyperparameters, one can get a quick and cheap evaluation of suitable hyperparameters. We mainly use two metrics to judge the performance of an ID head:
  • Accuracy of a $k$-NN classifier trained on a subset of the data (e.g. 10% of the data). This is relativley cheap to compute and can be done periodically during training. For the $k$-NN classifier, one can use either features of an encoder block or features of intermediate blocks in an ID head to judge the representation at the respective location in the network.
  • The accuracy of the NN-swap, i.e. how often is the NN from the NN-swap from the same class as the query sample. This metric is essentially free to compute as the NN-swap is required for training anyways.